### v1

# Discovery potential of a long-lived partner of inelastic dark matter at MATHUSLA in $U(1)_{X_{3}}$ extension of the standard model

Abstract
We investigate the discovery potential at the MATHUSLA experiment of a long-lived particle (LLP), which is the heavier state of inelastic scalar dark matter (DM) in third generation-philic $U(1)$ ($U(1)_{X_{3}}$) extension of the standard model. Since the heavier state and DM state form the complex scalar charged under the $U(1)_{X_{3}}$, it is natural that the heavier state $P$ is almost degenerate with the DM state and hence long-lived. We find that third generation-philic right-handed $U(1)$, $U(1)_{R_{3}}$, model is the most interesting, because third generation-philic models are less constrained by the current experimental results and right-handed $U(1)$ interactions leave visible final decay products without produing neutrinos. For a benchmark of the model parameters consistent with the current phenomenological constraints, we find that the travel distance of the LLP can be $\mathcal{O}(100)$ m and the LLP production cross section at the 14 TeV LHC can be $\mathcal{O}(10)$ fb. Thus, we conclude that the LLP can be discovered at the MATHUSLA with a sufficiently large number of LLP decay events inside the MATHUSLA detector.
preprint: EPHOU-24-014
I Introduction
Explanation of the neutrino oscillation phenomena between neutrino flavor species due to non-zero neutrino mass and existence of dark matter (DM) in the Universe require undiscovered particles and interactions of beyond the standard model (BSM) of particle physics. Nevertheless, any evidence of new particles in BSM have not yet been reported in any ongoing experiments. A reason of non-discovery of a new particle might be due to not its large mass but its extremely weak interaction strength, in other words, a relevant coupling constant is very small. Even if coupling constants and hence the resultant production cross sections of new particles are very small, we may expect certain amount of production of new particles if experiments are conducted with huge luminosity. The small coupling constant makes the new particle long-lived, and signals of such long-lived particles (LLPs), if they are electrically neutral, would be discovered through their “displaced vertex” signatures. In fact, displaced vertex signatures have been searched for at, for instance, ATLAS ATLAS:2022pib and FASER FASER:2023tle in Large Hadron Collider (LHC). Some near-future experiments such as FASER2 FASER:2018eoc will start soon, and other several future experiments have been proposed, targeting various range of LLP masses and decay lengths. Among such proposed future experiments for the displaced vertex search, MAssive Timing Hodoscope for Ultra Stable neutraL pArticles (MATHUSLA) experiment is striking, because it will be able to search LLPs with the decay length of the order of $100$Chou:2016lxi . Theoretical studies on search sensitivity for various BSM models have been examined in Refs. BhupalDev:2016nfr ; Evans:2017lvd ; Helo:2018qej ; Deppisch:2018eth ; Jana:2018rdf ; Bauer:2018uxu ; Curtin:2018ees ; Berlin:2018jbm ; Dercks:2018eua ; Deppisch:2019kvs ; No:2019gvl ; Wang:2019xvx ; Jodlowski:2019ycu ; Bolton:2019pcu ; Hirsch:2020klk ; Jana:2020qzn ; Gehrlein:2021hsk ; Sen:2021fha ; Guo:2021vpb ; Bhattacherjee:2021rml ; Du:2021cmt ; Kamada:2021cow ; Bertuzzo:2022ozu ; Liu:2022ugx ; Bandyopadhyay:2022mej ; Mao:2023zzk ; Jodlowski:2023yne ; Fitzpatrick:2023xks ; Curtin:2023skh ; Batz:2023zef ; Deppisch:2023sga ; Bernal:2023coo ; Bishara:2024rtp ; Liu:2024azc ; deVries:2024mla ; Liebersbach:2024kzc .
[…]
In the models of Ref. Okada:2019sbb , both DM state and the heavier state relevant for both co-annihilation and inelastic scattering of DM state originate from one complex scalar field. Since they belong to the single complex scalar field, those two states are naturally well degenerate. The decay rate of the heavier state into the lighter one (DM state) would be very small due to its very small phase space volume, so that the heavier state could be a natural candidate for the LLP that we are interested in. To examine the MATHUSLA prospects, inelastic DM models have been proposed Berlin:2018jbm ; Guo:2021vpb ; Bertuzzo:2022ozu  [^1] The mass splitting in Refs. Guo:2021vpb ; Bertuzzo:2022ozu is generated by an explicit gauge symmetry breaking term, and models have not been formulated in gauge invariant manner.. While the longevity of LLP in fermionic inelastic DM model proposed in Ref. Berlin:2018jbm is due to the mass degeneracy as well as its tiny kinetic mixing of dark photon, the LLP in our scalar inelastic DM model is long-lived due to only about $1\%$ mass degeneracy.
This paper is organized as follows: In Sec. II, we describe the extra gauge $U(1)$ interacting inelastic model, which is a variant of the original models in Ref. Okada:2019sbb . We identify a parameter set to reproduce thermal DM abundance and a benchmark point in Sec. III. In Sec. IV, for the benchmark point, we present a search prospect of LLPs by the MATHUSLA experiment. Section V is devoted to our summary.
II The Model
There are several possibilities of anomaly free gauged $U(1)$ extension of the standard model. Among them, the best studied model is the flavor universal $U(1)_{B-L}$ Pati:1973uk ; Davidson:1978pm ; Mohapatra:1980qe ; Mohapatra:1980 . However, because the new neutral gauge boson $Z^{\prime}$ in the $U(1)_{B-L}$ model interacts with the first and the second generation of SM fermions, the experimental constraints on the couplings are very stringent Carena:2004xs ; Amrith:2018yfb ; Das:2019fee . Thus, we may not expect a sizable production cross section for the LLP production from the $Z^{\prime}$ boson decay. Note that anomaly cancellation of $U(1)_{B-L}$ model is realized in each generation of fermions, hence it is generally possible for only a particular generation to be charged under the $U(1)$. The model where only the third generation fermions are charged, $U(1)_{(B-L)_{3}}$, is such a choice Okada:2019sbb ; Babu:2017olk ; Alonso:2017uky ; Bian:2017rpg ; Cox:2017rgn and well studied. As expected, the experimental constraints on the $U(1)_{(B-L)_{3}}$ model are much weaker compared with the universal $U(1)_{B-L}$ model delAmoSanchez:2010bt ; Faroughy:2016osc ; Aaboud:2017sjh ; Chun:2018ibr ; Elahi:2019drj . It is easy to generalize the $U(1)_{B-L}$ model to the $U(1)_{X}$ model by assigning the $U(1)$ charge for a field as a linear combination of its $U(1)_{Y}$ and $U(1)_{B-L}$ charge. Since $U(1)_{Y}$ and $U(1)_{B-L}$ are independently anomaly free, the $U(1)_{X}$ is automatically anomaly free Appelquist:2002mw ; Oda:2015gna ; Das:2016zue . A special case is so-called $U(1)_{R}$ model, where only right-handed (RH) SM fermions and RH neutrinos are charged under the $U(1)$ Jung:2009jz . In this paper, we generalize $U(1)_{(B-L)_{3}}$ to $U(1)_{X_{3}}$, where $X_{3}$ denotes a linear combination of $U(1)_{Y_{3}}$ and $U(1)_{(B-L)_{3}}$. We introduce an $U(1)_{X_{3}}$ charged scalar field $\phi_{1}$ with its charge $+1$ in addition to the minimal particle contents. The total particle contents are listed in Tab. 1. The field $\phi_{2}$ is responsible to break $U(1)_{X_{3}}$ gauge symmetry and $x_{H}$ is a real parameter which parameterize a combination weight of $(B-L)_{3}$ and $Y_{3}$.
SU(3)c SU(2)L U(1)Y U(1)${}_{X_{3}}$
$Q^{i}$ 3 2 $1/6$ $\left(\frac{1}{6}x_{H}+\frac{1}{3}\right)\delta_{i3}$
$u^{i}_{R}$ 3 1 $2/3$ $\left(\frac{2}{3}x_{H}+\frac{1}{3}\right)\delta_{i3}$
$d^{i}_{R}$ 3 1 $-1/3$ $\left(-\frac{1}{3}x_{H}+\frac{1}{3}\right)\delta_{i3}$
$L^{i}$ 1 2 $-1/2$ $\left(-\frac{1}{2}x_{H}-1\right)\delta_{i3}$
$e^{i}_{R}$ 1 1 $-1$ $\left(-x_{H}-1\right)\delta_{i3}$
$\Phi$ 1 2 $1/2$ $\frac{1}{2}x_{H}$
$N^{i}_{R}$ 1 1 00 $-\delta_{i3}$
$\phi_{1}$ 1 1 00 $+1$
$\phi_{2}$ 1 1 00 $+2$
Table 1: The particle contents of our $U(1)_{X_{3}}$ model. In addition to the SM particle content ($i=1,2,3$), three RH neutrinos ($N_{R}^{i}$ ($i=1,2,3$)) and two $U(1)_{X_{3}}$ scalar fields ($\phi_{1}$ and $\phi_{2}$) are introduced. The scalar $\phi_{2}$ is the Higgs field and develops its VEV to break the $U(1)_{X_{3}}$ symmetry, while $\phi_{1}$ has no VEV and includes the inelastic DM and its partner LLP as its components.
The $U(1)_{X_{3}}$ gauge interaction for an SM chiral fermion and RH neutrinos ($f_{L/R}$) can be read from the usual covariant derivative,
$\mathcal{L}_{\mathrm{int}}=\sum_{f_{j}}\overline{f_{j}}\gamma^{\mu}X_{\mu}g_{X_{3}}\left(q_{f_{jL}}P_{L}+q_{f_{jR}}P_{R}\right)f_{j},$ (1)
where $X^{\mu}$ is the $U(1)_{X_{3}}$ gauge field, $g_{X_{3}}$ is the gauge coupling constant, and $q_{f_{jL/R}}$ is a $U(1)_{X_{3}}$ charge of $f_{jL/R}$ (see Tab. 1).
The scalar potential, which is gauge-invariant and renormalizable, is expressed as Okada:2018xdh ; Chao:2017ilw
$V(\Phi,\phi_{1},\phi_{2})=$ $-M^{2}_{\Phi}|\Phi|^{2}+\frac{\lambda}{2}|\Phi|^{4}+M^{2}_{\phi_{1}}\phi_{1}\phi_{1}^{\dagger}-M^{2}_{\phi_{2}}\phi_{2}\phi_{2}^{\dagger}$
$+\frac{1}{2}\lambda_{1}(\phi_{1}\phi_{1}^{\dagger})^{2}+\frac{1}{2}\lambda_{2}(\phi_{2}\phi_{2}^{\dagger})^{2}+\lambda_{3}\phi_{1}\phi_{1}^{\dagger}(\phi_{2}\phi_{2}^{\dagger})$
$+(\lambda_{4}\phi_{1}\phi_{1}^{\dagger}+\lambda_{5}\phi_{2}\phi_{2}^{\dagger})|\Phi|^{2}-A(\phi_{1}\phi_{1}\phi_{2}^{\dagger}+\phi_{1}^{\dagger}\phi_{1}^{\dagger}\phi_{2}),$ (2)
with $\Phi$ being the SM Higgs doublet field. All parameters, $M^{2}_{i},\lambda_{i}$ and $A$, in the potential (2) are taken to be real and positive.
II.1 Dark matter mass and interactions
At the $U(1)_{X_{3}}$ and the electroweak (EW) symmetry breaking vacuum, the SM Higgs field and the $U(1)$ Higgs field are expanded around those VEVs, $v$ and $v_{2}$, as (in the unitary gauge)
$\Phi=$ $\left(\begin{array}[]{c}0\\ \frac{v+\varphi}{\sqrt{2}}\\ \end{array}\right),$ (5)
ϕ1=subscriptitalic-$\phi_{1}=$ $\frac{S+iP}{\sqrt{2}},$ (6)
ϕ2=subscriptitalic-$\phi_{2}=$ $\frac{v_{2}+\varphi_{2}}{\sqrt{2}}.$ (7)
The physical states ($\varphi$ and $\varphi_{2}$) are diagonalized to the mass eigenstates ($h$ and $H$) with masses $m_{h}$ and $m_{H}$ as
$\left(\begin{array}[]{c}\varphi\\ \varphi_{2}\\ \end{array}\right)=\left(\begin{array}[]{cc}\cos\alpha&\sin\alpha\\ -\sin\alpha&\cos\alpha\\ \end{array}\right)\left(\begin{array}[]{c}h\\ H\\ \end{array}\right).$ (8)
For a small mixing angle $\alpha$, $h$ is identified with the SM-like Higgs boson. Hereafter, we set $\alpha$ negligibly small. With the $U(1)_{X_{3}}$ and the EW symmetry breaking, the $Z^{\prime}$ boson which is the mass eigenstate after the $X$ becomes massive, $S$ and $P$ acquire their masses, respectively, as
$m_{Z^{\prime}}^{2}=$ $g_{X_{3}}^{2}4v_{2}^{2},$ (9)
$m_{S}^{2}=$ $M_{\phi_{1}}^{2}+\frac{1}{2}\lambda_{3}v_{2}^{2}+\frac{1}{2}\lambda_{4}v^{2}-\sqrt{2}Av_{2},$ (10)
$m_{P}^{2}=$ $M_{\phi_{1}}^{2}+\frac{1}{2}\lambda_{3}v_{2}^{2}+\frac{1}{2}\lambda_{4}v^{2}+\sqrt{2}Av_{2}.$ (11)
Note that the parameter $A$ controls the mass splitting between $S$ and $P$. Since we take $A$ positive, $S$ is lighter than $P$ and becomes the DM candidate. The other choice of $A<0$ causes no essential difference in our final results, except that $P$ is the DM particle in the case.
Gauge interaction of the DM particle is expressed as
$\mathcal{L}_{\mathrm{int}}=g_{X_{3}}Z^{\prime}{}^{\mu}\left((\partial_{\mu}S)P-S\partial_{\mu}P\right),$ (12)
and similarly 3rd generation quarks and leptons also interact with $Z^{\prime}$ boson with corresponding charges. The absence of $Z^{\prime}$-DM-DM coupling indicates that the $Z^{\prime}$-mediating DM scattering off with a nucleon is inelastic and ineffective for a mass splitting larger than the maximal energy transfer in the scatterings Hall:1997ah ; TuckerSmith:2001hy . Elastic scattering through Higgs bosons exchange Jungman:1995df can be neglected since we have set a very small Higgs mixing $\alpha$. We also set the coupling constant $A$ very small, so that $S$ and $P$ are well degenerate in mass. This degeneracy is crucial not only for the scalar $P$ being long-lived, but also for reproducing the observed DM relic density through $S$ and $P$ coannihilation process.
II.2 The decay width of $Z^{\prime}$ boson
The partial decay width of $Z^{\prime}\rightarrow f\bar{f}$ is given by
$\Gamma(Z^{\prime}\rightarrow f\bar{f})=$ $\frac{1}{48\pi m_{Z^{\prime}}^{2}}N_{c}|\mathcal{M}|^{2}\sqrt{m_{Z^{\prime}}^{2}-4m_{f}^{2}},$ (13)
$|\mathcal{M}|^{2}=$ $2g_{X_{3}}^{2}\left(q_{fL}^{2}\left(m_{Z^{\prime}}^{2}-m_{f}^{2}\right)+6q_{fL}q_{fR}m_{f}^{2}+q_{fR}^{2}\left(m_{Z^{\prime}}^{2}-m_{f}^{2}\right)\right),$ (14)
where the number of color $N_{c}$ is $3$ for quarks and $1$ for leptons. If $m_{f}\ll m_{Z^{\prime}}$, then Eq. (13) is reduced to
$\Gamma(Z^{\prime}\rightarrow f\bar{f})\simeq$ $\frac{g_{X_{3}}^{2}N_{c}m_{Z^{\prime}}}{24\pi}\left(q_{fL}^{2}+q_{fR}^{2}\right).$ (15)
We obtain the partial decay width of $Z^{\prime}$ into $S$ and $P$ as
$\Gamma(Z^{\prime}\rightarrow SP)=\frac{g_{X_{3}}^{2}}{48\pi}\frac{(m_{Z^{\prime}}^{2}-(m_{P}-m_{S})^{2})^{3/2}(m_{Z^{\prime}}^{2}-(m_{P}+m_{S})^{2})^{3/2}}{m_{Z^{\prime}}^{5}},$ (16)
from the vertex (12). The decay branching ratios are shown in Fig. 1. Note that $Br(Z^{\prime}\rightarrow\nu_{\tau}\nu_{\tau})$ vanishes at $x_{H}=-2$, which corresponds to $U(1)$ right-handed, $U(1)_{R}$. This fact plays an important role in the following discussion.
Figure 1: Branching ratios of the $Z^{\prime}$ decay for $g_{X}=0.1,m_{S}=130$ GeV and $m_{Z^{\prime}}=350$ GeV.
II.3 Decay of $P$
If $S$ and $P$ are strongly degenerate so that the mass difference is much smaller than the mass of $Z^{\prime}$, the two body decay of $P$ is kinematically forbidden. Thus, for the main decay mode, $P\rightarrow SZ^{\prime}{}^{*}\rightarrow Sf\bar{f}$, the total decay width ($\Gamma$), or equivalently the inverse of the lifetime ($\tau$),
[…]
is suppressed by the phase space volume Mohapatra:2023aei . As a result, $P$ is short-lived in cosmology but can be long-lived in collider experiments. The expected travel distance $c\tau$ for $x_{H}=-2$, in other words $U(1)_{R_{3}}$, is shown as the function of its mass $m_{P}$ in Fig. 2. We note that the mass difference,
$m_{P}^{2}-m_{S}^{2}=2\sqrt{2}Av_{2},$ (18)
from Eqs. (10) and (11), is always adjustable by suitably choosing the free parameter $A$ such that $c\tau$ becomes about $100$ m, which is ideal for the detection of $P$ decay inside the MATHUSLA detector. Here and hereafter, we have forcused on the $U(1)_{R_{3}}$ case. For $x_{H}\neq-2$, the decay mode of $P\rightarrow SZ^{\prime}{}^{*}\rightarrow S\nu_{\tau}\nu_{\tau}$ exits and becomes the dominant mode for $M_{P}-m_{S}\leq 2m_{\tau}$. In this case, the LPP $P$ provides only invisible decay products.
Figure 2: $c\tau$ for $g_{R_{3}}=0.1,m_{S}=150$ GeV and $m_{Z^{\prime}}=350$ GeV. Near the threshold $m_{P}-m_{S}\simeq 2m_{\tau}$ with $m_{\tau}$ being the tau-lepton mass, $c\tau$ is very long.
III Dark matter and LHC constraints
For evaluation of prospect at MATHUSLA in the next section, in this section, we will find a benchmark point which satisfies the thermal DM abundance and the latest LHC constraints. As metioned above, we concentrate on the $U(1)_{R_{3}}$ model. The LLP $P$ in other general $U(1)_{X_{3}}~{}(x_{H}\neq-2)$ may decay into neutrinos which are invisible for the MATHUSLA detector. Thus, the $U(1)_{R_{3}}$ model is the best choice from the viewpoint of the detection with a large decay length of $\mathcal{O}(100)$ m.
III.1 Thermal relic abundance
We estimate the thermal relic abundance of the real scalar DM, $S$, by solving the Boltzmann equation,
$\frac{dn}{dt}+3Hn=-\langle\sigma_{\mathrm{eff}}v\rangle(n^{2}-n_{\rm EQ}^{2}),$ (19)
where $H$ and $n_{\rm EQ}$ are the Hubble parameter and the DM number density at thermal equilibrium, respectively KolbTurner . In our model, the main annihilation mode is coannihilation $SP\rightarrow f\bar{f}$ through $s$-channel $Z^{\prime}$ exchange for $m_{Z^{\prime}}>m_{S}$ and the annihilation mode $SS\rightarrow Z^{\prime}Z^{\prime}$ by $u(t)$-channel $P$ exchange for $m_{Z^{\prime}}<m_{S}$ Okada:2019sbb . We use the effective thermal averaged annihilation cross section
$\langle\sigma_{\mathrm{eff}}v\rangle=\sum_{i,j=S,P}\langle\sigma_{ij}v_{ij}\rangle\frac{n_{i}}{n_{\rm EQ}}\frac{n_{j}}{n_{\rm EQ}},$ (20)
to include the coannihilation effects properly and $n$ in Eq. (19) should be understood as $n=\sum_{i}n_{i}$ for $i=S,P$ Griest:1990kh ; Edsjo:1997bg . The contours of thermal DM abundance $\Omega h^{2}\simeq 0.1$ Planck:2018vyg (blue curve) and $\mathrm{Br}(Z^{\prime}\rightarrow SP)$ (green curves) for the $U(1)_{R_{3}}$ model are shown in Fig. 3.
Figure 3: The contour (in blue) along which the observed DM relic abundance $\Omega h^{2}\simeq 0.1$ is reproduced for $g_{R_{3}}=0.1$ and $x_{H}=-2$.
III.2 Benchmark points
To find a viable benchmark point, we also take the $Z^{\prime}$ boson search at the LHC with ditau final state CMS:2024pjt ; ATLAS:2019erb into account. We evaluate the production cross section for the process $pp\rightarrow\tau^{+}\tau^{-}$ given by
$\sigma(pp\rightarrow Z^{\prime})\mathrm{BR}(Z^{\prime}\rightarrow\tau^{+}\tau^{-}),$ (21)
where the branching ratio is calculated from Eq. (13). As can be seen in Eq. (14), $\mathrm{BR}(Z^{\prime}\rightarrow\tau^{+}\tau^{-})$ depends on the value of $x_{H}$, and we forcus on $x_{H}=-2$ as stated above.
With the decay rates of $X$ estimated in Sec. II.2, as in Ref. Das:2019fee , since the total $X$ boson decay width is very narrow, we use the narrow width approximation to evaluate the $X$ boson production cross section
[…]
$\hat{\sigma}(\hat{s})$ $=\frac{4\pi^{2}}{3}\frac{\Gamma_{X}(X\rightarrow q\bar{q})}{m_{X}}\delta(\hat{s}-m_{X}^{2}),$ (23)
where $f_{q}$ and $f_{\bar{q}}$ are the parton distribution function (PDF) for a quark and antiquark ($b$ and $\bar{b}$ in our model), $\hat{s}=xys$ is the invariant mass squared of colliding quarks for the center of mass energy $s$. The factor $2$ in Eq. (22) counts two ways of $q$ coming from which proton out of two colliding protons. Since the most severe bound is from the dilepton channel ($\ell=\tau$), we calculate $\sigma(pp\rightarrow X)\mathrm{Br}(X\rightarrow\tau^{+}\tau^{-})$ and compare it with the CMS results CMS:2024pjt  [^2] The ATLAS bound ATLAS:2019erb is somewhat weaker than that from the CMS.. We employ PDFs of CTEQ6L Pumplin:2002vw with a factorization scale $Q=m_{Z^{\prime}}$ for simplicity.
The resultant production cross section is shown in the second column in Tab. 3. The production cross section is consistent with and smaller than the latest most stringent bound from the CMS CMS:2024pjt at the LHC. We summarize parameters and masses of relevant particles in Tab. 2.
$x_{H}$ $g_{X_{3}}$ $m_{Z^{\prime}}$ $m_{S}$
$BP_{-2}$ $-2$ $0.1$ $351$ GeV $150$ GeV
Table 2:
IV Prospect for MATHUSLA
The travel distance $c\tau\simeq 100$ m is the most sensitive range at MATHUSLA  Chou:2016lxi ; Curtin:2017izq ; Curtin:2017bxr . The pair production cross section of such a LLP $\chi$ for its $5\sigma$ discovery can be read as $\sigma_{\chi\chi}>0.3$ fb from Fig. 1 in Ref. Jana:2018rdf . On the other hand, in our scenario, the LLP $P$ is not pair produced but singly produced through $pp\rightarrow Z^{\prime}\rightarrow SP$. Thus, the required production cross section for the LLP $P$ to be discovered at the MATHUSLA would be set to be larger than $0.6$ fb. In Tab. 3, we list the $SP$ production cross section at $13$ TeV LHC and $14$ TeV LHC for our benchmark on Tab. 2 satisfying the DM abundance and the LHC bounds. We find that the MATHUSLA will be able to discover those LLPs because the production cross section $\sigma_{SP}$ is sufficiently larger than $0.6$ fb.
$\sigma(pp\rightarrow Z^{\prime}\rightarrow\tau\tau)$ (CMS bound on it) $\sigma_{SP}$$13$TeV LHC $\sigma_{SP}$$14$TeV LHC
$BP_{-2}$ $154$ fb ($<203$ fb) $10.8$ fb $12.9$ fb
Table 3:
V Summary
We have proposed a simple extension of the SM with an extra $U(1)_{R_{3}}$ gauge interaction with scalar particle of DM candidate with the charge $1$, third generation of SM fermions and third generation of right-handed neutrinos. After the SM singlet $U(1)_{R_{3}}$ breaking scalar with the charge $2$ develops the VEV, the gauge boson acquires the mass, and the tiny mass splitting between real and imaginary component of the charge $1$ scalar appears through the scalar tri-linear interaction. The lighter scalar $S$ is inelastic DM candidate with the heavier state $P$. Due to the mass degeneracy, the slightly heavier state $P$ is long-lived and will be able to be discovered as LLPs.
We have calculated the production cross section of LLP $P$ through the process $pp\rightarrow Z^{\prime}\rightarrow SP$ at the MATHUSLA for a benchmark point which satisfies stringent LHC bounds and thermal DM abundance. The production cross section of benchmark point turns out to be about $10$ fb, which is about an order of magnitude larger than the cross section required by $5\sigma$ discovery as a LLP at the MATHUSLA. Thus, we conclude that the heavier state of inelastic DM in our model can be discovered at the MATHUSLA.
Acknowledgments
This work is supported in part by the U.S. DOE Grant No. DE-SC0012447 and DE-SC0023713 (N.O.) and KAKENHI Grants No. JP23K03402 (O.S.).
Appendix A The decay rate of $P$
The spin averaged squared amplitude for $P(p)\rightarrow f(q_{1})\bar{f}(q_{2})S(q_{3})$ is given by
$\overline{|\mathcal{M}|^{2}}$
$=$ $\frac{2g_{R_{3}}^{4}}{m_{Z^{\prime}}^{4}\left(-2\left(p\cdot q_{3}\right)+m_{P}^{2}+m_{S}^{2}-m_{Z^{\prime}}^{2}\right)^{2}}\left(-2\left(-m_{P}^{2}+m_{S}^{2}+m_{Z^{\prime}}^{2}\right)^{2}\left(p\cdot q_{1}\right)^{2}\right.$
$-2\left(m_{P}^{2}-m_{S}^{2}+m_{Z^{\prime}}^{2}\right)^{2}\left(q_{1}\cdot q_{3}\right)^{2}+4\left(\left(m_{P}^{2}-m_{S}^{2}\right)^{2}-m_{Z^{\prime}}^{4}\right)\left(p\cdot q_{1}\right)\left(q_{1}\cdot q_{3}\right)$
$-2\left(-m_{P}^{2}+m_{S}^{2}+m_{Z^{\prime}}^{2}\right)^{2}\left(p\cdot q_{1}\right)\left(p\cdot q_{3}\right)+2\left(m_{P}^{2}-m_{S}^{2}+m_{Z^{\prime}}^{2}\right)^{2}\left(q_{1}\cdot q_{3}\right)\left(p\cdot q_{3}\right)$
$\left.+m_{f}^{2}\left(-2\left(\left(m_{P}^{2}-m_{S}^{2}\right)^{2}-m_{Z^{\prime}}^{4}\right)\left(p\cdot q_{3}\right)+m_{Z^{\prime}}^{4}\left(m_{P}^{2}+m_{S}^{2}\right)+\left(m_{P}^{2}-m_{S}^{2}\right)^{2}\left(m_{P}^{2}+m_{S}^{2}-2m_{Z^{\prime}}^{2}\right)\right)\right.$
$\left.+\left(p\cdot q_{1}\right)\left(m_{Z^{\prime}}^{4}\left(m_{P}^{2}-3m_{S}^{2}\right)+2m_{Z^{\prime}}^{2}\left(m_{S}^{4}-m_{P}^{4}\right)+\left(m_{P}^{2}-m_{S}^{2}\right)^{2}\left(m_{P}^{2}+m_{S}^{2}\right)\right)\right.$
$\left.+\left(q_{1}\cdot q_{3}\right)\left(m_{P}^{4}\left(m_{S}^{2}-2m_{Z^{\prime}}^{2}\right)+m_{P}^{2}\left(m_{S}^{4}+3m_{Z^{\prime}}^{4}\right)-m_{P}^{6}-\left(m_{S}^{3}-m_{S}m_{Z^{\prime}}^{2}\right)^{2}\right)\right).$ (24)
The integration of phase space volume is reduced to
$dQ=$ $\frac{d^{3}q_{1}}{(2\pi)^{3}2q_{1}^{0}}\frac{d^{3}q_{2}}{(2\pi)^{3}2q_{2}^{0}}\frac{d^{3}q_{3}}{(2\pi)^{3}2q_{3}^{0}}(2\pi)^{4}\delta^{(4)}\left(p-q_{1}-q_{2}-q_{3}\right)$
$=$ $\frac{1}{(2\pi)^{5}}\frac{d^{3}q_{1}}{2q_{1}^{0}}\frac{d^{3}q_{3}}{2q_{3}^{0}}\delta\left((m_{P}-q_{1}^{0}-q_{3}^{0})^{2}-(|\mathbf{q}_{1}|^{2}+|\mathbf{q}_{3}|^{2})-2|\mathbf{q}_{1}||\mathbf{q}_{3}|\cos\theta_{13}-m_{f}^{2}\right)$
$=$ $\frac{1}{8(2\pi)^{5}}dq_{1}^{0}dq_{3}^{0}d\Omega_{1}d\varphi_{13}d\cos\theta_{13}\delta\left(\frac{(m_{P}-q_{1}^{0}-q_{3}^{0})^{2}-|\mathbf{q}_{1}|^{2}-|\mathbf{q}_{3}|^{2}-m_{f}^{2}}{2|\mathbf{q}_{1}||\mathbf{q}_{3}|}-\cos\theta_{13}\right).$ (25)
Then, the integration range of energy turns out to be
$q_{3}^{\mathrm{min}}(q_{1}^{0})<q_{3}^{0}<q_{3}^{\mathrm{max}}(q_{1}^{0}),$ (26)
with
$q_{3}^{\mathrm{min}}(q_{1}^{0})=\frac{\left(m_{P}-q_{1}^{0}\right)\left(m_{P}^{2}-2m_{P}q_{1}^{0}+m_{S}^{2}\right)-\sqrt{\left(\left(q_{1}^{0}\right)^{2}-m_{f}^{2}\right)\left(\left(m_{P}^{2}-2m_{P}q_{1}^{0}-m_{S}^{2}\right)^{2}-4m_{f}^{2}m_{S}^{2}\right)}}{2\left(m_{f}^{2}+m_{P}\left(m_{P}-2q_{1}^{0}\right)\right)},$ (27)
$q_{3}^{\mathrm{max}}(q_{1}^{0})=\frac{\left(m_{P}-q_{1}^{0}\right)\left(m_{P}^{2}-2m_{P}q_{1}^{0}+m_{S}^{2}\right)+\sqrt{\left(\left(q_{1}^{0}\right)^{2}-m_{f}^{2}\right)\left(\left(m_{P}^{2}-2m_{P}q_{1}^{0}-m_{S}^{2}\right)^{2}-4m_{f}^{2}m_{S}^{2}\right)}}{2\left(m_{f}^{2}+m_{P}\left(m_{P}-2q_{1}^{0}\right)\right)},$ (28)
and
$m_{f}<q_{1}^{0}<\frac{m_{P}^{2}-m_{S}^{2}-2m_{f}m_{S}}{2m_{P}}.$ (29)
By integrating the differential decay rate
$d\Gamma=\frac{1}{2p^{0}}\overline{|\mathcal{M}|^{2}}dQ=\frac{1}{8(2\pi)^{3}m_{P}}\left[\overline{|\mathcal{M}|^{2}}\delta\left(\cdots-\cos\theta_{13}\right)d\cos\theta_{13}\right]dq_{1}^{0}dq_{3}^{0},$ (30)
for the range (26) and (29), we obtain the travel distance shown in Fig. 2.

### v2

@@
 Abstract
-We investigate the discovery potential at the MATHUSLA experiment of a long-lived particle (LLP), which is the heavier state of inelastic scalar dark matter (DM) in third generation-philic $U(1)$ ($U(1)_{X_{3}}$) extension of the standard model.
-Since the heavier state and DM state form the complex scalar charged under the $U(1)_{X_{3}}$, it is natural that the heavier state $P$ is almost degenerate with the DM state and hence long-lived.
-We find that third generation-philic right-handed $U(1)$, $U(1)_{R_{3}}$, model is the most interesting, because third generation-philic models are less constrained by the current experimental results and right-handed $U(1)$ interactions leave visible final decay products without produing neutrinos.
+We investigate the discovery potential at the MATHUSLA experiment of a long-lived particle (LLP), which is the heavier state of inelastic scalar dark matter (DM) in third-generation-philic $U(1)$ ($U(1)_{X_{3}}$) extension of the Standard Model.
+Since the heavier state and DM state form the complex scalar charged under the $U(1)_{X_{3}}$, it is natural that the heavier state $P$ is almost degenerate with the DM state and is hence long-lived.
+We find that third generation-philic right-handed $U(1)$, $U(1)_{R_{3}}$, model is the most interesting, because third-generation-philic models are less constrained by the current experimental results and right-handed $U(1)$ interactions leave visible final decay products without produing neutrinos.
 For a benchmark of the model parameters consistent with the current phenomenological constraints, we find that the travel distance of the LLP can be $\mathcal{O}(100)$ m and the LLP production cross section at the 14 TeV LHC can be $\mathcal{O}(10)$ fb.
@@
 I Introduction
-Explanation of the neutrino oscillation phenomena between neutrino flavor species due to non-zero neutrino mass and existence of dark matter (DM) in the Universe require undiscovered particles and interactions of beyond the standard model (BSM) of particle physics.
+Explanation of the neutrino oscillation phenomena between neutrino flavor species due to nonzero neutrino mass and the existence of dark matter (DM) in the Universe require undiscovered particles and interactions of beyond the standard model (BSM) of particle physics.
 Nevertheless, any evidence of new particles in BSM have not yet been reported in any ongoing experiments.
-A reason of non-discovery of a new particle might be due to not its large mass but its extremely weak interaction strength, in other words, a relevant coupling constant is very small.
-Even if coupling constants and hence the resultant production cross sections of new particles are very small, we may expect certain amount of production of new particles if experiments are conducted with huge luminosity.
+A reason of nondiscovery of a new particle might be due to not their large mass but their extremely weak interaction strength; in other words, the relevant coupling constant is very small.
+Even if coupling constants and hence the resultant production cross sections of new particles are very small, we may expect a certain amount of production of new particles if experiments are conducted with huge luminosity.
 The small coupling constant makes the new particle long-lived, and signals of such long-lived particles (LLPs), if they are electrically neutral, would be discovered through their “displaced vertex” signatures.
-In fact, displaced vertex signatures have been searched for at, for instance, ATLAS ATLAS:2022pib and FASER FASER:2023tle in Large Hadron Collider (LHC).
-Some near-future experiments such as FASER2 FASER:2018eoc will start soon, and other several future experiments have been proposed, targeting various range of LLP masses and decay lengths.
-Among such proposed future experiments for the displaced vertex search, MAssive Timing Hodoscope for Ultra Stable neutraL pArticles (MATHUSLA) experiment is striking, because it will be able to search LLPs with the decay length of the order of $100$Chou:2016lxi .
-Theoretical studies on search sensitivity for various BSM models have been examined in Refs.
+In fact, displaced vertex signatures have been sought at, for instance, ATLAS ATLAS:2022pib and FASER FASER:2023tle in the Large Hadron Collider (LHC).
+Some near-future experiments such as FASER2 FASER:2018eoc will start soon, and several other future experiments have been proposed, targeting various range of LLP masses and decay lengths.
+Among such proposed future experiments for the displaced vertex search, MAssive Timing Hodoscope for Ultra Stable neutraL pArticles (MATHUSLA) experiment is striking, because it will be able to search for LLPs with a decay length of the order of $100$Chou:2016lxi .
+Theoretical studies on the search sensitivity for various BSM models have been examined in Refs.
 BhupalDev:2016nfr ; Evans:2017lvd ; Helo:2018qej ; Deppisch:2018eth ; Jana:2018rdf ; Bauer:2018uxu ; Curtin:2018ees ; Berlin:2018jbm ; Dercks:2018eua ; Deppisch:2019kvs ; No:2019gvl ; Wang:2019xvx ; Jodlowski:2019ycu ; Bolton:2019pcu ; Hirsch:2020klk ; Jana:2020qzn ; Gehrlein:2021hsk ; Sen:2021fha ; Guo:2021vpb ; Bhattacherjee:2021rml ; Du:2021cmt ; Kamada:2021cow ; Bertuzzo:2022ozu ; Liu:2022ugx ; Bandyopadhyay:2022mej ; Mao:2023zzk ; Jodlowski:2023yne ; Fitzpatrick:2023xks ; Curtin:2023skh ; Batz:2023zef ; Deppisch:2023sga ; Bernal:2023coo ; Bishara:2024rtp ; Liu:2024azc ; deVries:2024mla ; Liebersbach:2024kzc .
@@
 In the models of Ref.
-Okada:2019sbb , both DM state and the heavier state relevant for both co-annihilation and inelastic scattering of DM state originate from one complex scalar field.
-Since they belong to the single complex scalar field, those two states are naturally well degenerate.
+Okada:2019sbb , both the DM state and the heavier state relevant for both coannihilation and inelastic scattering of the DM state originate from one complex scalar field.
+Since they belong to a single complex scalar field, those two states are naturally well degenerate.
 The decay rate of the heavier state into the lighter one (DM state) would be very small due to its very small phase space volume, so that the heavier state could be a natural candidate for the LLP that we are interested in.
@@
 While the longevity of LLP in fermionic inelastic DM model proposed in Ref.
-Berlin:2018jbm is due to the mass degeneracy as well as its tiny kinetic mixing of dark photon, the LLP in our scalar inelastic DM model is long-lived due to only about $1\%$ mass degeneracy.
-This paper is organized as follows: In Sec.
-II, we describe the extra gauge $U(1)$ interacting inelastic model, which is a variant of the original models in Ref.
+Berlin:2018jbm is due to the mass degeneracy as well as its tiny kinetic mixing of dark photons, the LLP in our scalar inelastic DM model is long-lived due to only about $1\%$ mass degeneracy.
+This paper is organized as follows: In Section II, we describe the extra gauge $U(1)$ interacting inelastic model, which is a variant of the original models in Ref.
 Okada:2019sbb .
-We identify a parameter set to reproduce thermal DM abundance and a benchmark point in Sec.
-III.
-In Sec.
-IV, for the benchmark point, we present a search prospect of LLPs by the MATHUSLA experiment.
+We identify a parameter set to reproduce thermal DM abundance and a benchmark point in Section III.
+In Section IV, for the benchmark point, we present the search prospects of LLPs by the MATHUSLA experiment.
 Section V is devoted to our summary.
 II The Model
-There are several possibilities of anomaly free gauged $U(1)$ extension of the standard model.
+There are several possibilities of the anomaly-free gauged $U(1)$ extension of the Standard Model.
 Among them, the best studied model is the flavor universal $U(1)_{B-L}$ Pati:1973uk ; Davidson:1978pm ; Mohapatra:1980qe ; Mohapatra:1980 .
@@
 Thus, we may not expect a sizable production cross section for the LLP production from the $Z^{\prime}$ boson decay.
-Note that anomaly cancellation of $U(1)_{B-L}$ model is realized in each generation of fermions, hence it is generally possible for only a particular generation to be charged under the $U(1)$.
-The model where only the third generation fermions are charged, $U(1)_{(B-L)_{3}}$, is such a choice Okada:2019sbb ; Babu:2017olk ; Alonso:2017uky ; Bian:2017rpg ; Cox:2017rgn and well studied.
+Note that anomaly cancellation of the $U(1)_{B-L}$ model is realized in each generation of fermions; hence it is generally possible for only a particular generation to be charged under $U(1)$.
+A model where only the third-generation fermions are charged, $U(1)_{(B-L)_{3}}$, is such a choice Okada:2019sbb ; Babu:2017olk ; Alonso:2017uky ; Bian:2017rpg ; Cox:2017rgn and has been well studied.
 As expected, the experimental constraints on the $U(1)_{(B-L)_{3}}$ model are much weaker compared with the universal $U(1)_{B-L}$ model delAmoSanchez:2010bt ; Faroughy:2016osc ; Aaboud:2017sjh ; Chun:2018ibr ; Elahi:2019drj .
-It is easy to generalize the $U(1)_{B-L}$ model to the $U(1)_{X}$ model by assigning the $U(1)$ charge for a field as a linear combination of its $U(1)_{Y}$ and $U(1)_{B-L}$ charge.
+It is easy to generalize the $U(1)_{B-L}$ model to the $U(1)_{X}$ model by assigning the $U(1)$ charge for a field as a linear combination of its $U(1)_{Y}$ and $U(1)_{B-L}$ charges.
 Since $U(1)_{Y}$ and $U(1)_{B-L}$ are independently anomaly free, the $U(1)_{X}$ is automatically anomaly free Appelquist:2002mw ; Oda:2015gna ; Das:2016zue .
-A special case is so-called $U(1)_{R}$ model, where only right-handed (RH) SM fermions and RH neutrinos are charged under the $U(1)$ Jung:2009jz .
+A special case is the so-called $U(1)_{R}$ model, where only right-handed (RH) SM fermions and RH neutrinos are charged under $U(1)$ Jung:2009jz .
 In this paper, we generalize $U(1)_{(B-L)_{3}}$ to $U(1)_{X_{3}}$, where $X_{3}$ denotes a linear combination of $U(1)_{Y_{3}}$ and $U(1)_{(B-L)_{3}}$.
 We introduce an $U(1)_{X_{3}}$ charged scalar field $\phi_{1}$ with its charge $+1$ in addition to the minimal particle contents.
-The total particle contents are listed in Tab.
-1.
-The field $\phi_{2}$ is responsible to break $U(1)_{X_{3}}$ gauge symmetry and $x_{H}$ is a real parameter which parameterize a combination weight of $(B-L)_{3}$ and $Y_{3}$.
+The total particle contents are listed in Table 1.
+The field $\phi_{2}$ is responsible for breaking the $U(1)_{X_{3}}$ gauge symmetry and $x_{H}$ is a real parameter that parameterizes a combination weight of $(B-L)_{3}$ and $Y_{3}$.
+Here, we note that the $U(1)_{(B-L)_{3}}$ and $U(1)_{X_{3}}$ models with a minimal Higgs doublet cannot reproduce realistic fermion masses and flavor mixings due to the gauge symmetry.
+Since the details of those UV completions are irrelevant for our main discussion on DM and LLP phenomenology, we work on the simplified particle content in Table 1.
+Nevertheless, we note an example of successful UV completions of $U(1)_{R_{3}}$ in Appendix A222 A few successful UV completions for $U(1)_{(B-L)_{3}}$ can be found in Refs.
+Babu:2017olk ; Alonso:2017uky ..
 SU(3)c SU(2)L U(1)Y U(1)${}_{X_{3}}$
@@
 $e^{i}_{R}$ 1 1 $-1$ $\left(-x_{H}-1\right)\delta_{i3}$
-$\Phi$ 1 2 $1/2$ $\frac{1}{2}x_{H}$
+$\Phi$ 1 2 $1/2$ 00
 $N^{i}_{R}$ 1 1 00 $-\delta_{i3}$
@@
 $\mathcal{L}_{\mathrm{int}}=\sum_{f_{j}}\overline{f_{j}}\gamma^{\mu}X_{\mu}g_{X_{3}}\left(q_{f_{jL}}P_{L}+q_{f_{jR}}P_{R}\right)f_{j},$ (1)
-where $X^{\mu}$ is the $U(1)_{X_{3}}$ gauge field, $g_{X_{3}}$ is the gauge coupling constant, and $q_{f_{jL/R}}$ is a $U(1)_{X_{3}}$ charge of $f_{jL/R}$ (see Tab. 1).
-The scalar potential, which is gauge-invariant and renormalizable, is expressed as Okada:2018xdh ; Chao:2017ilw
+where $X^{\mu}$ is the $U(1)_{X_{3}}$ gauge field, $g_{X_{3}}$ is the gauge coupling constant, and $q_{f_{jL/R}}$ is a $U(1)_{X_{3}}$ charge of $f_{jL/R}$ (see Table 1).
+Here, we do not consider the gauge kinetic mixing because the mixing between $Z$ and $Z^{\prime}$ bosons is experimentally severely constrained; in fact the mixing can also be canceled in a UV model (see Appendix A.).
+According to the experimental constraints, we have chosen our renormalization condition so as to vanish the kinetic mixing at the scale of $P$ mass.
+The scalar potential, which is gauge invariant and renormalizable, is expressed as Okada:2018xdh ; Chao:2017ilw
 $V(\Phi,\phi_{1},\phi_{2})=$ $-M^{2}_{\Phi}|\Phi|^{2}+\frac{\lambda}{2}|\Phi|^{4}+M^{2}_{\phi_{1}}\phi_{1}\phi_{1}^{\dagger}-M^{2}_{\phi_{2}}\phi_{2}\phi_{2}^{\dagger}$
@@
 II.1 Dark matter mass and interactions
-At the $U(1)_{X_{3}}$ and the electroweak (EW) symmetry breaking vacuum, the SM Higgs field and the $U(1)$ Higgs field are expanded around those VEVs, $v$ and $v_{2}$, as (in the unitary gauge)
+At $U(1)_{X_{3}}$ and the electroweak (EW) symmetry breaking vacuum, the SM Higgs field and the $U(1)$ Higgs field are expanded around the VEVs $v$ and $v_{2}$ as (in the unitary gauge)
 $\Phi=$ $\left(\begin{array}[]{c}0\\ \frac{v+\varphi}{\sqrt{2}}\\ \end{array}\right),$ (5)
@@
 Hereafter, we set $\alpha$ negligibly small.
-With the $U(1)_{X_{3}}$ and the EW symmetry breaking, the $Z^{\prime}$ boson which is the mass eigenstate after the $X$ becomes massive, $S$ and $P$ acquire their masses, respectively, as
+With $U(1)_{X_{3}}$ and the EW symmetry breaking, the $Z^{\prime}$ boson (the mass eigenstate after the $X$ becomes massive), $S$ and $P$ acquire their masses, respectively, as
 $m_{Z^{\prime}}^{2}=$ $g_{X_{3}}^{2}4v_{2}^{2},$ (9)
@@
 $\mathcal{L}_{\mathrm{int}}=g_{X_{3}}Z^{\prime}{}^{\mu}\left((\partial_{\mu}S)P-S\partial_{\mu}P\right),$ (12)
-and similarly 3rd generation quarks and leptons also interact with $Z^{\prime}$ boson with corresponding charges.
-The absence of $Z^{\prime}$-DM-DM coupling indicates that the $Z^{\prime}$-mediating DM scattering off with a nucleon is inelastic and ineffective for a mass splitting larger than the maximal energy transfer in the scatterings Hall:1997ah ; TuckerSmith:2001hy .
-Elastic scattering through Higgs bosons exchange Jungman:1995df can be neglected since we have set a very small Higgs mixing $\alpha$.
+and similarly third-generation quarks and leptons also interact with the $Z^{\prime}$ boson with corresponding charges.
+The absence of $Z^{\prime}$-DM-DM coupling indicates that the $Z^{\prime}$-mediating DM scattering with a nucleon is inelastic and ineffective for a mass splitting larger than the maximal energy transfer in the scatterings Hall:1997ah ; TuckerSmith:2001hy .
+Elastic scattering through Higgs bosons exchange Jungman:1995df can be negligible for a very small Higgs mixing $\alpha$ and the coupling $\lambda_{4}$.
 We also set the coupling constant $A$ very small, so that $S$ and $P$ are well degenerate in mass.
@@
 $|\mathcal{M}|^{2}=$ $2g_{X_{3}}^{2}\left(q_{fL}^{2}\left(m_{Z^{\prime}}^{2}-m_{f}^{2}\right)+6q_{fL}q_{fR}m_{f}^{2}+q_{fR}^{2}\left(m_{Z^{\prime}}^{2}-m_{f}^{2}\right)\right),$ (14)
-where the number of color $N_{c}$ is $3$ for quarks and $1$ for leptons. If $m_{f}\ll m_{Z^{\prime}}$, then Eq. (13) is reduced to
+where the color number $N_{c}$ is $3$ for quarks and $1$ for leptons. If $m_{f}\ll m_{Z^{\prime}}$, then Eq. (13) is reduced to
 $\Gamma(Z^{\prime}\rightarrow f\bar{f})\simeq$ $\frac{g_{X_{3}}^{2}N_{c}m_{Z^{\prime}}}{24\pi}\left(q_{fL}^{2}+q_{fR}^{2}\right).$ (15)
@@
 1.
-Note that $Br(Z^{\prime}\rightarrow\nu_{\tau}\nu_{\tau})$ vanishes at $x_{H}=-2$, which corresponds to $U(1)$ right-handed, $U(1)_{R}$.
+Note that $Br(Z^{\prime}\rightarrow\nu_{\tau}\nu_{\tau})$ vanishes at $x_{H}=-2$, which corresponds to right-handed $U(1)$, $U(1)_{R}$.
 This fact plays an important role in the following discussion.
@@
 II.3 Decay of $P$
-If $S$ and $P$ are strongly degenerate so that the mass difference is much smaller than the mass of $Z^{\prime}$, the two body decay of $P$ is kinematically forbidden.
+If $S$ and $P$ are strongly degenerate so that the mass difference is much smaller than the mass of $Z^{\prime}$, the two-body decay of $P$ is kinematically forbidden.
 Thus, for the main decay mode, $P\rightarrow SZ^{\prime}{}^{*}\rightarrow Sf\bar{f}$, the total decay width ($\Gamma$), or equivalently the inverse of the lifetime ($\tau$),
@@
 III Dark matter and LHC constraints
-For evaluation of prospect at MATHUSLA in the next section, in this section, we will find a benchmark point which satisfies the thermal DM abundance and the latest LHC constraints.
+For the evaluation of prospects at MATHUSLA in the next section, in this section, we will find a benchmark point thst satisfies the thermal DM abundance and the latest LHC constraints.
 As metioned above, we concentrate on the $U(1)_{R_{3}}$ model.
@@
 Figure 3: The contour (in blue) along which the observed DM relic abundance $\Omega h^{2}\simeq 0.1$ is reproduced for $g_{R_{3}}=0.1$ and $x_{H}=-2$.
+Green curves are contours of $\mathrm{Br}(Z^{\prime}\rightarrow SP)$.
+Here, $g_{R_{3}}=0.1$ in the left pannel and $=0.2$ in the right pannel.
 III.2 Benchmark points
-To find a viable benchmark point, we also take the $Z^{\prime}$ boson search at the LHC with ditau final state CMS:2024pjt ; ATLAS:2019erb into account.
-We evaluate the production cross section for the process $pp\rightarrow\tau^{+}\tau^{-}$ given by
-$\sigma(pp\rightarrow Z^{\prime})\mathrm{BR}(Z^{\prime}\rightarrow\tau^{+}\tau^{-}),$ (21)
-where the branching ratio is calculated from Eq.
-(13).
-As can be seen in Eq.
-(14), $\mathrm{BR}(Z^{\prime}\rightarrow\tau^{+}\tau^{-})$ depends on the value of $x_{H}$, and we forcus on $x_{H}=-2$ as stated above.
-With the decay rates of $X$ estimated in Sec.
-II.2, as in Ref.
-Das:2019fee , since the total $X$ boson decay width is very narrow, we use the narrow width approximation to evaluate the $X$ boson production cross section
+To find a viable benchmark point, we take the $Z^{\prime}$ boson search at the LHC into account.
+With the decay rates of $Z^{\prime}$ estimated in Section II.2, since the total $Z^{\prime}$ boson decay width is very narrow, as in Ref.
+Das:2019fee , we use the narrow width approximation to evaluate the $Z^{\prime}$ boson production cross section
 […]
@@
 The factor $2$ in Eq.
-(22) counts two ways of $q$ coming from which proton out of two colliding protons.
-Since the most severe bound is from the dilepton channel ($\ell=\tau$), we calculate $\sigma(pp\rightarrow X)\mathrm{Br}(X\rightarrow\tau^{+}\tau^{-})$ and compare it with the CMS results CMS:2024pjt  [^2] The ATLAS bound ATLAS:2019erb is somewhat weaker than that from the CMS..
+(21) accounts for two ways of $q$ coming from which proton out of two colliding protons.
 We employ PDFs of CTEQ6L Pumplin:2002vw with a factorization scale $Q=m_{Z^{\prime}}$ for simplicity.
-The resultant production cross section is shown in the second column in Tab.
-3.
+Since the severest bound on the $Z^{\prime}$ boson is from the dilepton channel ($\ell=\tau$CMS:2024pjt ; ATLAS:2019erb 333 We note that the present experimental bound from the $Z^{\prime}\rightarrow Zh$ mode reported by the ATLAS ATLAS:2022enb and CMS CMS:2024phk Collaborations for $m_{Z^{\prime}}=\mathcal{O}(100)$ GeV are about one order of magnitude weaker than that from the $\tau^{+}\tau^{-}$ mode., we evaluate the production cross section for the process $pp\rightarrow\tau^{+}\tau^{-}$ given by
+$\sigma(pp\rightarrow Z^{\prime})\mathrm{BR}(Z^{\prime}\rightarrow\tau^{+}\tau^{-}),$ (23)
+where the branching ratio is calculated from Eq.
+(13), and compare it with the CMS results CMS:2024pjt  [^4] The ATLAS bound ATLAS:2019erb is somewhat weaker than that from the CMS..
+The resultant production cross section is shown in the second column in Table 3.
 The production cross section is consistent with and smaller than the latest most stringent bound from the CMS CMS:2024pjt at the LHC.
-We summarize parameters and masses of relevant particles in Tab.
-2.
+We summarize parameters and masses of relevant particles of benchmark points with different values of the $U(1)_{R_{3}}$ gauge coupling constant in Table 2.
 $x_{H}$ $g_{X_{3}}$ $m_{Z^{\prime}}$ $m_{S}$
-$BP_{-2}$ $-2$ $0.1$ $351$ GeV $150$ GeV
-Table 2:
+$BP_{1}$ $-2$ $0.1$ $351$ GeV $150$ GeV
+$BP_{2}$ $-2$ $0.2$ $850$ GeV $351$ GeV
+Table 2: The values of parameters and the masses of $Z^{\prime}$ and $S$ in benchmark points.
 IV Prospect for MATHUSLA
 The travel distance $c\tau\simeq 100$ m is the most sensitive range at MATHUSLA  Chou:2016lxi ; Curtin:2017izq ; Curtin:2017bxr .
-The pair production cross section of such a LLP $\chi$ for its $5\sigma$ discovery can be read as $\sigma_{\chi\chi}>0.3$ fb from Fig.
+The pair production cross section of such an LLP $\chi$ for its $5\sigma$ discovery can be read as $\sigma_{\chi\chi}>0.3$ fb from Fig.
 1 in Ref.
@@
 On the other hand, in our scenario, the LLP $P$ is not pair produced but singly produced through $pp\rightarrow Z^{\prime}\rightarrow SP$.
-Thus, the required production cross section for the LLP $P$ to be discovered at the MATHUSLA would be set to be larger than $0.6$ fb.
-In Tab.
-3, we list the $SP$ production cross section at $13$ TeV LHC and $14$ TeV LHC for our benchmark on Tab.
-2 satisfying the DM abundance and the LHC bounds.
-We find that the MATHUSLA will be able to discover those LLPs because the production cross section $\sigma_{SP}$ is sufficiently larger than $0.6$ fb.
+Thus, the required production cross section for the LLP $P$ to be discovered at MATHUSLA would be set to be larger than $0.6$ fb.
+In Table 3, we list the $SP$ production cross section at $13$ TeV LHC and $14$ TeV LHC for our benchmark on Table 2 satisfying the DM abundance and the LHC bounds.
+We find that MATHUSLA will be able to discover those LLPs because the production cross section $\sigma_{SP}$ is sufficiently larger than $0.6$ fb.
 $\sigma(pp\rightarrow Z^{\prime}\rightarrow\tau\tau)$ (CMS bound on it) $\sigma_{SP}$$13$TeV LHC $\sigma_{SP}$$14$TeV LHC
-$BP_{-2}$ $154$ fb ($<203$ fb) $10.8$ fb $12.9$ fb
-Table 3:
+$BP_{1}$ $154$ fb ($<203$ fb) $10.8$ fb $12.9$ fb
+$BP_{2}$ $7.78$ fb ($<8.24$ fb) $0.697$ fb $0.888$ fb
+Table 3: Estimated production cross section for tau pairs and $SP$ pairs in benchmark points.
 V Summary
-We have proposed a simple extension of the SM with an extra $U(1)_{R_{3}}$ gauge interaction with scalar particle of DM candidate with the charge $1$, third generation of SM fermions and third generation of right-handed neutrinos.
-After the SM singlet $U(1)_{R_{3}}$ breaking scalar with the charge $2$ develops the VEV, the gauge boson acquires the mass, and the tiny mass splitting between real and imaginary component of the charge $1$ scalar appears through the scalar tri-linear interaction.
-The lighter scalar $S$ is inelastic DM candidate with the heavier state $P$.
+We have proposed a simple extension of the SM with an extra $U(1)_{R_{3}}$ gauge interaction with a scalar particle of DM candidate with the charge $1$, third generation of SM fermions and third generation of right-handed neutrinos.
+After the SM singlet $U(1)_{R_{3}}$ breaking scalar with charge $2$ develops a VEV, the gauge boson acquires a mass, and a tiny mass splitting between the real and imaginary components of the charge $1$ scalar appears through the scalar trilinear interaction.
+The lighter scalar $S$ is an inelastic DM candidate with a heavier state $P$.
 Due to the mass degeneracy, the slightly heavier state $P$ is long-lived and will be able to be discovered as LLPs.
-We have calculated the production cross section of LLP $P$ through the process $pp\rightarrow Z^{\prime}\rightarrow SP$ at the MATHUSLA for a benchmark point which satisfies stringent LHC bounds and thermal DM abundance.
-The production cross section of benchmark point turns out to be about $10$ fb, which is about an order of magnitude larger than the cross section required by $5\sigma$ discovery as a LLP at the MATHUSLA.
-Thus, we conclude that the heavier state of inelastic DM in our model can be discovered at the MATHUSLA.
+We have calculated the production cross section of LLP $P$ through the process $pp\rightarrow Z^{\prime}\rightarrow SP$ at MATHUSLA for a benchmark point that satisfies stringent LHC bounds and thermal DM abundance.
+The production cross section of the benchmark point turns out to be about $10$ fb, which is about an order of magnitude larger than the cross section required by $5\sigma$ discovery as a LLP at the MATHUSLA.
+Thus, we conclude that the heavier state of inelastic DM in our model can be discovered at MATHUSLA.
 Acknowledgments
 This work is supported in part by the U.S. DOE Grant No. DE-SC0012447 and DE-SC0023713 (N.O.) and KAKENHI Grants No. JP23K03402 (O.S.).
-Appendix A The decay rate of $P$
+Appendix A An extension of Higgs sector
+We note a possible UV completed model to generate top qurak mass and how gauge kinetic mixing parameter affects the mixing between $Z$ and $Z^{\prime}$.
+We may consider an extension with three Higgs doublet fields $\Phi_{1},\Phi_{2}$ and $\Phi_{3}$ and some singlet scalar fields to eliminate unwanted NG modes.
+The charge assignments of those Higgs fields are in Tab.
+4.
+Those doublet fields develop the VEVs as
+$\Phi_{i}=$ $\left(\begin{array}[]{c}0\\ \frac{v_{i}}{\sqrt{2}}\\ \end{array}\right).$ (26)
+The quark Yukawa interactions are expressed as
+$\mathcal{L}_{\mathrm{Yukawa}}=$ $\sum_{i,j}^{1,2}\left(-y_{ij}^{u}\overline{Q}_{i}\tilde{\Phi}_{1}u_{jR}-y_{ij}^{d}\overline{Q}_{i}\Phi_{1}d_{jR}-y_{3j}^{u}\overline{Q}_{3}\tilde{\Phi}_{1}u_{jR}-y_{3j}^{d}\overline{Q}_{3}\Phi_{1}d_{jR}\right)$
+$-y_{i3}^{u}\overline{Q}_{i}\tilde{\Phi}_{2}u_{3R}-y_{i3}^{d}\overline{Q}_{i}\Phi_{2}d_{3R}-y_{33}^{u}\overline{Q}_{3}\tilde{\Phi}_{2}u_{3R}-y_{33}^{d}\overline{Q}_{3}\Phi_{2}d_{3R}+\mathrm{H.c.},$ (27)
+and there are non-vanishing entries for all $3$ by $3$ elements.
+The mass of top quark is generated through the VEV $v_{2}$.
+At the same time, the mass matrix of neutral gauge bosons $\hat{Z}$ and $\hat{X}$ in the interaction basis are given by
+$\left(\begin{array}[]{cc}M_{\hat{Z}\hat{Z}}^{2}&M_{\hat{Z}\hat{X}}^{2}\\ M_{\hat{X}\hat{Z}}^{2}&M_{\hat{X}\hat{X}}^{2}\\ \end{array}\right)=\left(\begin{array}[]{cc}\frac{g_{2}^{2}v^{2}}{8c_{W}^{2}}&-\frac{g_{2}g_{X}}{4c_{W}}\left(-v_{2}^{2}+q_{X_{3}}v_{3}^{2}\right)\\ -\frac{g_{2}g_{X}}{4c_{W}}\left(-v_{2}^{2}+q_{X_{3}}v_{3}^{2}\right)&\frac{g_{X}^{2}}{2}\left(v_{2}^{2}+q_{X_{3}}^{2}v_{3}^{2}+4v_{\phi}^{2}\right)\\ \end{array}\right),$ (28)
+with $v^{2}=v_{1}^{2}+v_{2}^{2}+v_{3}^{2}$, where $g_{2}$ is the $SU(2)_{L}$ gauge coupling, $g_{X}$ is the $U(1)_{R_{3}}$ gauge coupling, $v_{\phi}$ is the VEV of singlet Higgs field with the $U(1)_{R_{3}}$ charge $-2$, the the $U(1)_{R_{3}}$ charges of $\Phi_{1},\Phi_{2}$ and $\Phi_{3}$ are taken to be $0,-1,q_{X_{3}}$, respectively and $c_{W}$ is the cosine of the Weinberg angle.
+Note that the mass mixing between $\hat{Z}$ and $\hat{X}$ is tiny for $-v_{2}^{2}+q_{X_{3}}v_{3}^{2}\simeq 0$ Seto:2020jal .
+The field redefinition by the orthogonal matrix resolve the gauge kinetic mixing
+$\mathcal{L}_{\mathrm{Gauge}}\supset\frac{\sin\epsilon}{2}\hat{B}_{\mu\nu}\hat{X}^{\mu\nu},$ (29)
+but induces additional mass mixings. The resultant mass matrix for the canonical kinetic terms becomes
+$\left(\begin{array}[]{cc}M^{2}_{\hat{Z}\hat{Z}}&\frac{M^{2}_{\hat{Z}\hat{X}}}{c_{\epsilon}}-M^{2}_{\hat{Z}\hat{Z}}s_{W}t_{\epsilon}\\ \frac{M^{2}_{\hat{Z}\hat{X}}}{c_{\epsilon}}-M^{2}_{\hat{Z}\hat{Z}}s_{W}t_{\epsilon}&\frac{s_{W}c_{\epsilon}t_{\epsilon}}{c_{\epsilon}^{2}}\left(M^{2}_{\hat{Z}\hat{Z}}s_{W}c_{\epsilon}t_{\epsilon}-2M^{2}_{\hat{Z}\hat{X}}\right)+\frac{1}{c_{\epsilon}^{2}}M^{2}_{\hat{X}\hat{X}}\\ \end{array}\right).$ (30)
+The magnitude of the off-diagonal component, or equvalently the $Z-Z^{\prime}$ mixing, is by experiments stringently constrained as $M^{2}_{\hat{Z}\hat{X}}-M^{2}_{\hat{Z}\hat{Z}}s_{W}s_{\epsilon}\lesssim(0.1\,\mathrm{GeV})^{2}$ ParticleDataGroup:2024cfk , so that the shift of the $Z$ boson mass due to the mixing must be within the error of its measurement.
+In our model, by seting the ratio $v_{2}/v_{3}$ in $M^{2}_{\hat{Z}\hat{X}}$ appropriately so that the off-diagonal element can be as small as to be consistent, the mixing between $Z$ and $Z^{\prime}$ is always negligible even under non-vanishing gauge kinetic parameter $\epsilon$.
+One can easily see that physics of dark matter and $Z^{\prime}$ described in the main text are intact for the above extension.
+SU(3)c SU(2)L U(1)Y U(1)${}_{R_{3}}$ $Z_{2}$
+$Q^{i}$ 3 2 $1/6$ 00 $+$
+$u^{i}_{R}$ 3 1 $2/3$ $-\delta_{i3}$ $+$
+$d^{i}_{R}$ 3 1 $-1/3$ $\delta_{i3}$ $+$
+$\Phi_{1}$ 1 2 $1/2$ 00 $+$
+$\Phi_{2}$ 1 2 $1/2$ $-1$ $+$
+$\Phi_{3}$ 1 2 $1/2$ $q_{X3}$ $+$
+$\phi_{1}$ 1 1 00 $+1$-
+$\phi_{1}^{\prime}$ 1 1 00 $+1$ $+$
+$\phi_{2}$ 1 1 00 $+2$ $+$
+$\phi_{3}$ 1 1 00 $q_{X3}$ $+$
+Table 4: The short list of particle contents with an extended Higgs sector.
+$\phi_{1}^{\prime}$ and $\phi_{3}$ provide the masses of NG bosons to be heavy psuedo scalars.
+A $Z_{2}$ parity distinguishes $\phi_{1}$ and $\phi_{1}^{\prime}$, and let DM stable.
+Appendix B The decay rate of $P$
 The spin averaged squared amplitude for $P(p)\rightarrow f(q_{1})\bar{f}(q_{2})S(q_{3})$ is given by